MEMORANDUM

To: J. Albertine, D. Helms, W. Skinner

cc: Division (M7), FEL Coordination Group

From: F. Dylla

Subject: IRFEL Weekly Report, August 25-29, 1997

Date: August 29, 1997

Management

This was a busy and productive week for FEL installation and sub-system commissioning activities: the linac cryomodule was successfully cooled and remains stable at 2K; the completed injector dump beamline was put under a warm nitrogen purge prior to pumpdown next week; the initial check-out of the Personnel Safety System (PSS) was successfully completed; the first arc stands were installed; and the recirculation dump was installed.

Final revisions were collected on the first set of IR FEL proposal summaries that will be forwarded next week to the DOE-BES program office.

A meeting was held on Wednesday of the ARC University Coordinating committee to discuss the final lay-out of the ARC Building laboratories. We were able to co-locate many similar activities between the four universities and Jefferson Lab, particularly laser/optics labs and materials analysis labs that will have obvious benefits to FEL users, FEL development and Jefferson Lab R&D activities. The ARC universities are making plans for moving into the ARC Building at the end of the calendar year.

We confirmed the attendance at next month's SURA/DOE Science and Technology Review (Sept.17-19) of Steve Laderman, from Hewlett Packard Research Laboratories. Steve will be reviewing both the FEL program and the laboratory's relations with industry. In his position at Hewlett Packard, Steve oversees the company's interests at the Stanford Synchrotron Radiation Laboratory.

Accelerator Systems

Regarding the beam-transport system:

Optical Chicane Dipoles - The eight optical chicane dipoles continued their measurements on the measurement stand with the results showing that five are acceptable if the calibration of the stand is correct. Three exhibited unacceptable increase in the field integral at one edge of the 9 cm required good field region. On that side, the field integral rises to 2.5 parts per ten thousand at 100 amps (Though the rise in integral is not seen at 200 amps - the one-additional-cryomodule-upgrade current). We went through a round of flipping the position of the field clamp plates end for end on one of these outlyers. The results first indicated that the magnet joined the population of good magnets with later measurements indicating that it reverted back to its original increase in the field integral at one edge. Suspicious that the variation is an artifact of the measurement apparatus, we are ending the week with the a full set of measurements with the entire magnet flipped end for end in the apparatus. Interpretation of the results awaits Jeff Karn's return on Tuesday. As soon as we achieve a greater level of confidence in the measurements and have a handle on correcting the outlyers, we will pin the field clamps in their final position and authorize installation of a group of four. We will use one of the remaining four to use as a second reference dipole in a series of tests to obtain absolute calibration of the standard dipole and the data set.

Other Dipoles - All the gluing of the Injection/Extraction DU/DV style dipoles' mu metal/brass shims was completed this week at Everson Electric and the first magnets of each were shipped. The core slabs for the 180 degree dipoles are being pinned together. They are slated for delivery early next week. The first of the 180 Degree dipole coil sets was potted this week. The bids for assembly of these dipoles are due at COB Tuesday, so we will know where to ship the coils, cores and parts coincident with their availability. The remaining parts of the 180 degree magnet measurement apparatus are out for bid, requesting a late September delivery.

In the quadrupole and sextupole area, winding of coils for the eight Panofsky trim quads continued at Magnet Enterprises International (MEI). The three quadrupole magnets for the energy-recovery dump line (QIs) continue in final assembly at New England Technicoil with delivery delayed to early next week because of come inconsistencies in the size of the specified coil positioning fillers.

In the corrector area, manufacture of the coils for the phasing dipoles (DGs) was delayed at MEI because of the late delivery of rectangular conductor. Manufacture of the mu metal cores for phasing dipoles (DGs) and reverse bend correctors (DF&DCs) is still unsettled, awaiting bids from the mu metal supplier. Coils for the reverse bend correctors are out for bids due at COB on Tuesday.

Drawings for the air core corrector coil sub assemblies have gone through a second back check with refinements resulting from experience in building the prototype DB. The drawings are due for signature on Wednesday. The prototype horizontal air core corrector was completed and the magnet mounted in the test stand. We intend to run the pivotal cross talk test that duplicates the intended proximity to nearby quadrupole's iron by drawing sign-off time. Passing the cross-talk test is needed before approval. Design of the skew-quadrupole magnets remains in final detailing.

Some optical chicane chambers are going through a second cycle of cleaning (which will require an off-line bake) after large particulates were found in one. The particle counter test employed earlier detects small particulates only and didn't detect this class of contamination. This incident redoubles our vigilance during tube cleaning.

Chamber welding - Inspection revealed problems that required scrapping two bodies of the five straight reverse bend chambers. Parts for recovery are due next week as the welding manpower for these critical elements was increased to three qualified welders. The remaining three bodies are machined to accept flanges. The seven curved reverse bend chambers are at 15 % completion using improved procedures. The parts for chambers for the 180 degree dipoles are now complete at Master Machine and they are developing weld procedures on samples using argon hydrogen mixture for shield gas, a significant advantage for welding this thicker joint.

The first articles for stands for the arcs arrived this week and are installed.

In the power supply area, the two power supplies for injection and extraction dipoles have been shipped by air freight with customs inspection requiring an additional week before we receive them. The remaining probes for zeroing the dipoles at the dumps were received and the electronics for the systems are on hand. Brackets to mount the probes remain at the concept level but are considered to be easily created.

In the installation area, the entire beam line in the injector to its dump region is under vacuum and robust. The line downstream of the cryomodule and into the recirculation dump is now receiving assembly attention.

Regarding cryomodule fabrication

Cryomodule

The installed cryomodule is cooled down and filled with 2 K helium. The cryomodule is operating at 2 K and will start preliminary check out next week. Full RF commissioning will begin when RF and safety systems are operational.

Warm Windows

The warm window program continues to progress on three fronts. The primary program, modification of a cold window to the warm location, has had a successful resonance ring test of a completed window to 50 kW cw RF power. The heating was limited to 90 degrees celsius on the ceramic and 45 degrees celsius on the flange. These temperatures are well within requirements. This window suffers from a vacuum leak that is not related to RF power operations. Additional windows are in fabrication at this time and are not expected to leak. The remaining two efforts are making progress toward completion of prototype assemblies. Prototypes should be available for evaluation during September.

Regarding RF systems:

The elbow for the 500 kV High Voltage Power Supply was installed for pressure and high voltage testing. We ran atmospheric pressure high voltage tests on the and were able to hold 150-160 kV. This meets our requirements, but without a lot of margin. The Cathode Power Supply wiring continues for the required PSS interlocks of the Buncher system. The RF system for the Buncher is about 50 % complete. The installation of the 50 kW klystron systems for the quarter cryomodule is about 50 % complete. Hipotronics technicians made the necessary repairs to the first 225 kW Variable Power Supply. AC power will be connected next week to allow high voltage testing. Hipotronics will send people back to repair the second unit after the first unit has been high voltage tested. The RF system for the linac is about 40 % complete. The Master Oscillator (M.O.) is ready for phase noise testing with the drive laser. Insulation for the LCW pipe and M.O. distribution cables were received today. The M.O. system is about 80 % complete. The SOP for operation of the 50 kW RF systems is complete and ready for use.

Regarding instrumentation and control systems:

All of the beam viewers, BPM's and multislit from the gun to the injector dump are installed. The installation of peripherals will commence post bake-out. The 100mm camera lenses are beginning to arrive, all of the CCD cameras are here and being checked out. Talks are ongoing with Princeton Plasma Physics Laboratory to borrow their polychromator to analyze bunch rms bunch length of the FEL electron beam.

The PSS certification went well, a short punch list is being worked, the wiring and remote interfaces are being finished up this week.

The instrumentation for the dump water skid is being assembled, this will be run in early September.

The vacuum system control cable pull is complete. These are being terminated now. The software controls checkout will start next week.

The AMS backplane board is due back from fabrication next week. One potential hold up is the delay in shipping of one of the card-edge connectors, we are working this. The production AMS buffer board prototype is complete and being evaluated, this should be ready for production next week. The analog-to-fiber chassis arrived, the initial tests are disappointing, the frequency response falls short of 1 MHz. This unit is used to transmit the AMS signals to the MCC (~.5 km). Internal to the FEL building the AMS is flat to over 20 MHz.

The DC power trim racks are being powered up and checked out. After the bake-out of the injector is complete, the magnets will be connected and field polarity checked.

FEL Systems

Optics

Work continued this week on the drive laser and the optical system. The laser's delivery optics are complete to the Brewster window in the Clean Room. A He-Ne laser collinear with the drive laser beam is used to align the transport optics on the ground floor. We discovered two of the stands were designed for the wrong elevation, these have been modified. The first mirror can is hung and aligned, we are proceeding with alignment through the telescope box. The two cans being built by the Machine Shop are being cleaned before welding. The drawing for the mirror cassettes and stands are signed and in Procurement. We are nearing completion on the design for the insertable dumps. We discovered that the molybdenum mirrors for the optical collimator were figured incorrectly, they are going back to the vendor for repolishing. We will temporarily use mirrors dielectrically coated for 633 nm. We received the 5.5" dia. window needed in the upstream optical cavity assembly, the entire assembly is nearly complete. Michelle Shinn had a video conference with staff in the physics branch at China Lake. This group does the metrology on our optics. She answered questions about the specifications for the IR Demo FEL, and participated in a discussion on how to quickly get better optics. Of the several action items that came out of that meeting, those for Jefferson Lab personnel await the return of staff on travel at the International FEL Conference. She also learned that a test set of mirrors made by Rocky Mountain Instrument met specifications for reflectivity. They will begin measurements on the scatter next week, that value, combined with the reflectivity will give us a value for the coating absorption.

Commissioning/Operations

The second meeting of FEL principal investigators was held and along with discussion of PI duties and responsibilities, a list of personnel assigned to each task for the commissioning was generated.

The Commissioning meeting discussed progress of the magnet performance integration plan and then used the present DW magnet measurements as an application/test of the process. The DW measurements indicate integrated field flatness errors on three of the seven magnets measured at 2.5 to 3 times the level of the specification. From the integration plan perspective, the magnets allowed us to use the criteria stated for when a magnet is good enough; i.e. can it be modeled? is it reproducible? etc. They provided a good first test of how magnets that fall outside the specification should be handled.

The injector dump setup procedure was completed and modeling is continuing on the buncher and cryounit fine setup procedure.

Facility

The majority of activities in the facility this week focused on installation items noted above in the system reports. We continue to improve the housekeeping in the accelerator vault as the vacuum system activities are ramped up. Floor areas are being cleaned, all non-necessary installation material is being removed, and additional mats are being laid in front of all building entrances. A meeting was held with the contractor to resolve the improper leveling of the floor near some of the perimeter drains which will help floor cleaning exercises when resolved. The contractor continues to deal with equipment needed to certify the elevator. All other punch list items are minor.